OpenStudy (anonymous):

WILL GIVE MEDAL: A website that sells songs for downloading increased its price per song from $0.99 to $1.29. Ariana spent $15.36 downloading songs before and after the price increase. She downloaded 4 more songs at $0.99 than at $1.29. The set of equations below represents the situation where x is the number of songs Ariana downloaded at $0.99 and y is the number of songs she downloaded at $1.29. x = y + 4 .99x + 1.29y = 15.36 What is the exact number of songs Ariana downloaded at the $0.99 price?

OpenStudy (anonymous):

@jdoe0001 @ganeshie8 please help... I'm desperate

OpenStudy (jdoe0001):

so.. if you were to solve say \(\bf x = y + 4\) for "y"... what would that give you?

OpenStudy (anonymous):

4x = y ?

OpenStudy (jdoe0001):

hmm you may want to brush up your linear equation simplification

OpenStudy (anonymous):

@jdoe0001 could you help me with how to solve x= y + 4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

can you please help me i need to turn this in today

OpenStudy (jdoe0001):

well... anyone writing it down, so you can transcribe it over, isn't going to "help" you though you need to do linear simplification on that, then substitute on the 2nd equation and then simplify the combined equation so.... you need to brush up your linear simplification bit firstly -> http://www.youtube.com/watch?v=n3FbQGGKVfc

OpenStudy (anonymous):

@jdoe0001 does y = 8

OpenStudy (anonymous):

@jdoe0001 are you still going to help me??

OpenStudy (jdoe0001):

well.... I think you were helped already..... you need to cover the linear simplication section first

OpenStudy (jdoe0001):

as I said... anyone just writing it down so you can "transcribe it" over, is not really "help

OpenStudy (anonymous):

i watched the video you linked, and as a final result for the x=y+4 equation i got y=8

OpenStudy (anonymous):

@jdoe0001